### v1

# A Survey on Tabular Data Generation: Utility, Alignment, Fidelity, Privacy, and Beyond

Abstract
Generative modelling has become the standard approach for synthesising tabular data. However, different use cases demand synthetic data to comply with different requirements to be useful in practice. In this survey, we review deep generative modelling approaches for tabular data from the perspective of four types of requirements: utility of the synthetic data, alignment of the synthetic data with domain-specific knowledge, statistical fidelity of the synthetic data distribution compared to the real data distribution, and privacy-preserving capabilities. We group the approaches along two levels of granularity: (i) based on the primary type of requirements they address and (ii) according to the underlying model they utilise. Additionally, we summarise the appropriate evaluation methods for each requirement and the specific characteristics of each model type. Finally, we discuss future directions for the field, along with opportunities to improve the current evaluation methods. Overall, this survey can be seen as a user guide to tabular data generation: helping readers navigate available models and evaluation methods to find those best suited to their needs.
Model
Primary
Requirement
Model
Type
Mixed
Data
Maximum
# Features
Generation
Time
Utility
Alignment
Fidelity
Privacy
CWGAN [?] Utility GAN 36
OCT-GAN [?] GAN 59 medium
GANBLR [?] GAN & BN 55
TabGFN [?] GAN & Flow Networks 37
TabDDPM [?] Diffusion 51 slow
STaSy [?] Diffusion 58 slow
CoDi [?] Diffusion 31 medium
GReaT [?] LLM 47 slow
TabPFGen [?] Energy-based model 77
ITS-GAN [?] Alignment GAN & AE 45
C3-TGAN [?] GAN & BN 54
GOGGLE [?] VAE & GNN 168 medium
TDGGD [?] Diffusion 44
C-DGM [?] Neural Network Layer 109 fast ✓✓
DGM+DRL [?] Neural Network Layer 64 fast ✓✓
TGAN [?] Fidelity GAN 55 medium
CTGAN [?] GAN 785 fast
CTAB-GAN [?] GAN 55 medium
CTAB-GAN+ [?] GAN 55 medium
TVAE [?] VAE 785 fast
Neural Spline Flows [?] Normalising Flows 50
FinDiff [?] Diffusion 84 slow
AutoDiff [?] Diffusion & AE 61 slow
TabSyn [?] Diffusion & VAE 48 medium
medGAN [?] Privacy GAN & AE 1071 fast
IT-GAN [?] GAN & AE 59 slow
TableGAN [?] GAN 32 medium
PATE-GAN [?] GAN 617 fast
CuTS [?] GAN 20
Table 1: Overview of the surveyed methods proposed for synthesising tabular data. For each work, we report (i) the requirement type it primarily focuses on, (ii) the model type, (iii) whether the model allows for mixed data types (marked with “✓” if so, and with “✗” otherwise), (iv) the maximum number of features used for evaluation (indicating whether the model can handle high-dimensionality data), (v) the model’s data generation time relative to the other models. In the last four columns, we indicate with “✓” (resp., “✗”) the requirements that were (resp., were not) addressed, specifically, (vi) utility, (vii) alignment, (viii) fidelity, and (ix) privacy. Note that “—” indicates that data are not available, and “✓✓” indicates that the method provides guarantees that the domain-specific constraints are satisfied.
1 Introduction
In recent years, the synthesis of realistic tabular data has emerged as a critical research area, driven by the need for privacy-preserving machine learning, robust AI benchmarking, and data augmentation in high-stakes applications such as finance and healthcare. These diverse needs led to not only a proliferation of different generative models, but also to an explosion of diverse evaluation metrics whose aim is to evaluate whether a model is able to address all the requirements of the given use case. However, the sheer diversity of these metrics has made it increasingly difficult to assess and compare different synthesisers in a standardised manner.
This survey takes a unique stance and analyses models and evaluations methods alike from the point of view of the user, categorising them on the ground of the requirements they primarily address and measure: utility, alignment, fidelity, and privacy. The analysis thus serves as a structured resource for both newcomers and experienced practitioners seeking a formal evaluation framework for tabular data synthesis. To this end, we first give an overview of the tabular data synthesis problem and its requirements (§2). Then, we examine in detail existing methods along two dimensions, grouping them (i) by the requirement they primarily focus on and (ii) by the model architecture type they are based on (§3-6). For each requirement, we detail the relevant evaluation protocols and metrics. A summary of the approaches surveyed in these is given in Table 1, which also gives an overview of the characteristics for each model, such as their ability to handle high-dimensional and mixed data types, as well as their sample generation times. We lastly review other deep generative models originally developed for other fields (such as natural language processing and computer vision), which can be adapted to tabular data synthesis (§7). We conclude with a discussion on related surveys and future directions and opportunities for improvement (§8).
2 Synthesising Tabular Data
Tabular Dataset.
In the tabular data synthesis field, a tabular dataset $\mathcal{D}$ can be usually defined as a triple $\mathcal{D}=(\mathcal{\mathcal{X},\mathcal{A},\mathcal{V}})$ where: (i) $\mathcal{X}=\{{x}^{1},\ldots,{x}^{N}\}$ is a finite set of instances (or rows), with each $\mathbf{x}^{i}$ (with $i=1,\ldots,N$) representing an individual data point, (ii) $\mathcal{A}=\{A^{1},\ldots,A^{K}\}$ is a finite set of attributes (or columns), each representing a feature of the data points, and (iii) $\mathcal{V}:\mathcal{X}\times\mathcal{A}\to\mathbb{D}$ is a value assignment function, which maps every data-point-feature pair $({x}^{i},A^{j})$ into an element of the domain $\mathbb{D}^{j}$, i.e., the set of permissible values for the attribute $A^{j}$. One of the many challenges of this field is given by the fact that the domain of each feature can be very different from one another. Common domains are: binary domains, i.e., $\mathbb{D}=\{0,1\}$, categorical domains, e.g., $\mathbb{D}=\{\text{Red},\text{Yellow},\text{Blue}\}$, and numerical domains, i.e., $\mathbb{D}\subseteq\mathbb{R}$. Given a dataset $\mathcal{D}$ with $N>1$, it is always possible to partition its instances in order to obtain a training set used to train a machine learning model, and a test set used to test a machine learning model.
Deep Generative Modelling for Tabular Data. To formulate the problem of tabular data generation, the literature makes the assumptions that given a tabular dataset $\mathcal{D}$, there exists an unknown distribution $p_{X}$ over the random variables $X\in\bigcup_{j=1}^{K}\mathbb{D}^{j}$ from which $\mathcal{D}$ was created by drawing $N$ i.i.d. samples. The goal of standard generative modeling is to learn from $\mathcal{D}$ the parameters $\theta$ of a generative model such that the model distribution $p_{\theta}$ approximates $p_{X}$. While synthesising tabular data is a long-standing problem [?], in this survey, we focus on methods that use deep generative models, i.e., models based on deep neural networks.
Requirements over Synthetic Data. While above we have the goal for standard generative modelling, we know that different use cases can lead to additional requirements, which lead to different goals for the generative task. In this survey, we review the following requirements:
1. Utility requirement: Synthetic data should yield similar predictive performance to real data when used to train machine learning (ML) models for the same task, such as classification or regression. For instance, given a healthcare dataset like WiDS [?], which contains medical records of patients from the first 24 hours of intensive care, and the target column specifies whether a patient has been diagnosed with diabetes mellitus, training a classifier on synthetic data should yield comparable (or better) performance on the real test set to that obtained by training the classifier on real data of the same size. Hence, given an ML model $m$ and a predictive performance metric $r$ (e.g., accuracy), maximising utility entails learning from $\mathcal{D}$ the parameters $\theta$ of a generative model such that training $m$ on the dataset $\mathcal{D}^{\prime}$ sampled from $p_{\theta}$ leads to the best score according to $r$.
2. Alignment requirement: Synthetic data should align with any known (user-provided) domain-specific knowledge. The knowledge can be expressed in multiple ways: simply as constraints on the range of values that a single feature can assume, or as more complex constraints capturing relations between the features. Continuing the example above, if the minimum and the maximum haemoglobin level are two of the features recorded for each patient, then clearly the real data do not contain any records for which the value of the minimum is higher than the value of the maximum level. This type of constraint is easily captured using a linear inequality. Hence, given a set of constraints expressing some background knowledge about the sample space of $p_{X}$, i.e., stating which samples are admissible and which are not, the goal is to learn the parameters $\theta$ of a generative model such that (i) the model distribution $p_{\theta}$ approximates $p_{X}$, and (ii) the sample space of $p_{\theta}$ is compliant with the constraints.
3. Fidelity requirement. Synthetic data should preserve the statistical properties of the real data. Continuing with our example, if we are maximising fidelity on the WiDS dataset, then the distribution of the synthetic values of the minimum haemoglobin level feature should resemble the real distribution of the same feature. Hence, in this case, the goal is to learn parameters $\theta$ such that the marginals of $p_{\theta}$ are as similar as possible to the corresponding marginals of $p_{X}$, and the joint distribution $p_{\theta}$ should closely follow $p_{X}$.
4. Privacy requirement. Synthetic data should present minimal risk of disclosing sensitive attributes and of re-identifying individuals from the real data. Continuing the earlier example, the WiDS dataset contains sensitive features such as the patient’s age, whether the patient had an elective surgery, and whether the patient received a leukemia diagnosis. If a rare combination of values for these features is retained in the synthetic data, then a patient might be identified, hence failing to meet privacy requirements. Hence, in this case, the goal is to learn the parameters $\theta$ such that it is impossible to sample a data point from $p_{\theta}$ that would allow for the re-identification (or for disclosing sensitive attributes) of a real data point.
In what follows, we group existing deep generative modelling methods for synthesising tabular data by the requirement they primarily address.
Figure 1: Visualisation of common model types used for synthesising tabular data. Here, $x$, $z$, and $\tilde{x}$ denote a real sample, noise sample, and synthetic sample, respectively; $\tilde{x}^{\prime}$ is a sample modified by layer-based models; $x_{i}$ is the sample at step $i$ of forward diffusion (for $i$ in $\{1,\ldots,T\}$, where $T$ is the maximum number of diffusion steps); and $\tilde{x}_{i}$ is the sample at step $i$ of reverse diffusion.
3 Utility
Determining whether synthetic data can replace real training data in downstream tasks has been the most common evaluation approach (e.g., [????]). The main reason for this is that synthetic data are often generated to augment or create datasets for real-world applications, where good predictive performance is crucial, e.g., predicting diabetes diagnoses from ICU data [?]. Therefore, utility still remains a key requirement for synthetic data and benchmark for comparing synthesisers.
3.1 Evaluation Protocol and Metrics
The most common protocol used to evaluate the utility of generative models is the Train on Synthetic, Test on Real (TSTR) Protocol (see, e.g., [?]). The procedure is to train a suite of ML models (e.g., a support vector machine, XGBoost) on the synthetic data, and then test their performance on a real test set, reporting the performance using different metrics on the ground of the task defined by the dataset of interest: for classification (resp., regression) datasets, standard metrics include accuracy and F1 score (resp., root mean square error and explained variance).
3.2 Methods
GAN-based models. Many of the works surveyed here are based on the original Generative Adversarial Network (GAN) [?], which relies on two neural networks: (i) a generator, which given a noise vector returns a synthetic datapoint and hence learns the real data distribution, and (ii) a discriminator, which given a data point (either real or synthetic) classifies it as either real or synthetic and hence learns to distinguish between synthetic and real data. The GAN jointly trains these two models in an adversarial framework, where the generator’s objective is to produce samples that confuse the discriminator. See the first schema from the left in Figure 1 for a simple abstraction of GAN-based models. CWGAN: Building on the seminal Wasserstein GAN (WGAN) [?], which used the Wasserstein distance as a loss function for a GAN model, ? [?] proposed CWGAN as an oversampling framework for class balancing. CWGAN generates synthetic samples for underrepresented classes, augmenting the training set. Results indicate that oversampling is beneficial primarily for strongly non-linear datasets. OCT-GAN: proposed by ? [?], incorporates neural ordinary differential equations (neural ODEs) in their GAN through the introduction of a new layer that allows for the extraction of a sequence of hidden vector representations given an input sample, i.e., the hidden evolution trajectory of the sample. The trajectory is used to help the discriminator decide whether a sample is real or synthetic. While improving utility, the usage of ODEs makes OCT-GAN slower than most GAN-based models.
GAN and BN-based models. GANBLR: ? (?) construct the generator and discriminator as Bayesian networks (BNs), which are able to encode known feature interactions. The background knowledge inclusion leads to better utility than standard GANs.
GAN and flow-networks-based models. TabGFN: ? (?) (i) model conditional distributions between the features using flow networks [?], where the features are ordered and sequentially generated, making it possible to trace the generation process and construct a sampling process trajectory; and (ii) adopt the WGAN-GP [?] critic with gradient penalty as a reward function to learn these relationships and the order of features. A key limitation is that TabGFN cannot handle continuous features, which can affect performance on high-dimensional continuous data having large variance.
Diffusion-based models. TabDDPM: Motivated by the their success in computer vision, ? ? introduced TabDDPM: a diffusion-based model for synthesising tabular data. Popular for its high utility results, TabDDPM is able to model both discrete and continuous features, but separately, by using multinomial [?] and Gaussian diffusion models [?], respectively. For a simple abstraction on how a diffusion model can be used to generate data, see the central schema in Figure 1. STaSy: Proposed by ? [?], STaSy is based on a score-based generative model (SGM), which was shown to be a competitive alternative to diffusion-based models in [?]. Contrarily to TabDDPM, StaSy estimates the score function (gradient of log-probability) instead of explicitly modelling the forward and reverse Markov processes. As the loss is difficult to train, STaSy uses a self-paced learning strategy, which starts with a subset of training data—yielding a stable, low loss—and gradually expands to the full dataset. Evaluated on 15 real-world datasets against 7 baselines (mostly GAN-based models), STaSy achieves high utility. Due to the forward and backward passes, both StaSy and TabDDPM suffer from slow sampling times. Moreover, the base SGM model is known to be unstable in high-dimensional settings. CoDi: Motivated by the difficulty of modelling discrete features, ? [?] proposed CoDi, a framework comprising co-evolving diffusion models that separately handle continuous and discrete features as in TabDDPM [?], with the difference that the two models condition on each other during training. At each training step, the discrete (resp., continuous) diffusion model takes the perturbed sample from the continuous (resp., discrete) one as input, and both models are conditioned on both samples from the previous step during the denoising process. Slower than most GAN-based models (excluding OCT-GAN), CoDi is faster than STaSy and TabDDPM, while also obtaining higher utility on mixed types datasets.
LLM-based models. GReaT [?] represents an answer to the abundance of works that convert tabular data into numerical representations thus losing the contextual connections between the features, which very often carry semantic meaning. Indeed, GReaT follows these steps (also illustrated in Figure 1): first, the tabular data undergoes a textual encoding process, converting it into text. Next, a feature order permutation step is applied. The resulting sentences are then used to fine-tune the large language models (LLMs) [?]. At sampling time, either a single feature name or arbitrary feature-value pairs are given as input to the LLM, which then completes them (note that this allows for arbitrary conditioning). Using this approach, GReaT outperforms most of the baselines w.r.t. utility.
Energy-based models. TabPFGen [?] is a generative model that uses a pretrained network as an energy-based model for data augmentation and class balancing. Although its capabilities are limited to low-dimensional inputs (it was tested on datasets with maximum 10 features), TabPFGen demonstrates good utility performance in the data augmentation task by simply using a pretrained model compared to specialised GAN-, VAE-, and diffusion-based models.
4 Alignment
Brought to the attention of the research community more recently [?] than the other requirements, background knowledge alignment is an important condition for synthetic data when evaluating its realism.
4.1 Evaluation Protocol and Metrics
In [?], three metrics have been proposed to evaluate alignment: (i) the Constraint Violation Rate (CVR), which computes the percentage of samples that do not satisfy at least one of the constraints in the available set of constraints, (ii) Constraint Violation Coverage (CVC), computing the percentage of constraints that have been violated at least once by the sample set, and (iii) the sample-wise Constraint Violation Coverage (sCVC) which determines the average percentage of samples violating each of the constraints. Out of these three metrics, CVR has also been used in [?], albeit called feasibility rate there.
4.2 Methods
GAN and AE-based models. ITS-GAN: proposed for tabular data augmentation ITS-GAN [?] is an early method advocating for the use of background knowledge. The knowledge it can express is of two types: (i) rules stating that the value of one feature uniquely determines the value of another (e.g., the feature Position determines the feature Salary), and (ii) rules stating that a specific value for a set of features determines the specific value for another (e.g., if Position = “CEO” then Salary = “2M”). To encode the first types of rules, they train on the real data an autoencoder (AE) for each rule to output the value of the dependent features given the values of the independent ones. Then, at training time, the GAN generator is trained to minimise the difference between its output and one of the AEs, while also being penalised if it violates any rule of the second type. Additionally, the discriminator receives as input the difference between the AE’s outputs and the synthetic samples, thus being able to use this information to assess the sample.
GAN and BN-based models. C3-TGAN: Similarly to GANBLR, C3-TGAN [?] uses Bayesian networks to capture background knowledge. However, while GANBLR models only discrete features, C3-TGAN can handle both discrete and continuous features by following the CTGAN [?] approach. To incorporate explicit attribute correlations and property constraints from background knowledge, it represents constraints as control vectors (similar to the conditional vectors in CTGAN) and uses them to guide training, ensuring better alignment.
VAE and GNN-based models. GOGGLE: [?] is a different approach which integrates knowledge about pairwise feature dependencies by encoding them in a graph where each feature is a node and an edge exists if a relation between the two features is known. New relations can be learnt using a message passing mechanism from graph neural networks (GNNs), which is jointly trained with a VAE-based architecture, where only relationships prioritised by the learned graph influence feature generation. Due to its relational learning graph-based component, it might be difficult to scale GOGGLE to datasets of higher dimensionality.
Diffusion-based models. TDGGD: The first framework encoding background knowledge about feature relations into diffusion-based models is TDGGD [?]. The relations can express lower and upper bounds over single features or their sum. Rather than explicitly modelling the relations, TDGGD models whether the features are part of such relations or not, which only gives an indication of hidden relations between the columns to the model.
Neural network layer-based models. C-DGM: The first method to constrain deep generative models and guarantee the constraints satisfaction was proposed in [?]. The constraints can capture any set of linear inequalities over the feature space. Differently from the methods surveyed above, this approach relies on a layer to be added right before the sample output layer, which restricts the output space of the model to coincide with the space defined by the constraints. The layer can be added on top of any deep generative model (DGM), and the resulting models are called C-DGMs. For a visual representation on how the layer-based methods can be added on a GAN model, see the right-most schema in Figure 1. As the layer is differentiable and acyclic, it can be added both at inference and training time, while also improving the utility of the model. DGM+DRL: The above layer was further extended in  [?], in order to capture constraints as expressive as disjunctions over linear inequalities (which allow for expressing relations like “if the sum of two features is greater than 10 then the difference of other two features should be lower than 5”), thus modelling non-convex and even disconnected output spaces. The new layer is called Disjunctive Refinement Layer (DRL), and a DGM constructed using DRL is called DGM+DRL. Just as C-DGM, DGM+DRL is able to guarantee alignment with the background knowledge, while also improving utility across all considered baselines.
5 Fidelity
While statistical fidelity is not a good predictor of how well the synthetic data would perform in downstream tasks [?], it can provide useful observations on how the synthetic data compares to the real data. Fidelity was indeed one of the requirements often measured in early works on synthesising tabular data.
5.1 Evaluation Protocol and Metrics
Being the longest-standing requirements, a large number of methods were developed to evaluate the statistical fidelity of synthetic data w.r.t. real data. [^1] In this survey, we focus only on quantitative evaluations. These can be categorised into (i) feature-wise, (ii) pair-wise, and (iii) joint evaluations.
Feature-wise evaluation: compares synthetic and real data feature by feature, using different metrics based on feature type. For continuous features, Wasserstein distance and the Kolmogorov-Smirnov test are common, while Jensen-Shannon divergence or total variation distance are common metrics for discrete features. The difference in treatment is supported empirically in [?], which found that Jensen-Shannon divergence is unstable for measuring the statistical similarity between continuous features’ distributions, especially with no overlap between synthetic and real data.
Pair-wise evaluation: investigates how well relationships between pairs of features are preserved in the synthetic w.r.t. the real data. Common metrics include: (i) RMSE (and MAE) differences in mutual information between pairs of features, and (ii) correlation difference, using Pearson’s coefficient for continuous pairs, Theil’s uncertainty coefficient or the total variation distance for discrete pairs, and the correlation ratio between discrete-continuous pairs of features.
Joint evaluation: compares the joint distribution of the synthetic samples against the real samples’ distribution. It is more difficult to determine how similar are two (potentially high-dimensional) joint distributions. Thus, such an analysis is often conducted in a controlled environment, e.g., on simulated data, rather than real-world data. Relevant evaluation methods include measuring the likelihood fitness score, separately for real and synthetic data (as in [?]).
5.2 Methods
GAN-based models. TGAN [?] is an early GAN-based model that proposes using long short-term memory networks with an attention mechanism to synthesise data sequentially, feature by feature, in order to model tabular data of mixed types. Comparing with conventional statistical models and showing better fidelity performance, TGAN contributed to the popularity of the GAN-based approaches for tabular data. CTGAN: focused on better modelling the types of the variables in an effort to increase fidelity. Indeed, the authors propose a new preprocessing method, which is column-type-specific and models discrete features in the continuous space via Gumbel-Softmax transformations, while it employs mode-specific normalisation via a variational Gaussian mixture model applied for each continuous features. Thanks to these transformations, CTGAN is able to avoid mode collapse, which often burdens other GAN-based models, like medGAN [?] and TableGAN [?]. CTAB-GAN: Building upon CTGAN, CTAB-GAN [?] enhances statistical fidelity by addressing data imbalance and long-tail issues. It introduces a conditional vector to handle mixed-type features and multi-modal continuous variables, extending support beyond discrete and continuous features. Further, it supports mixed-type features: which are features that might have highly-recurring discrete values, but also continuous values across the data points (e.g., a feature Credit Card Transaction USD might often have value 00, but it can also exceed $500$). The benefit of providing support for mixed-type features is reflected in the fidelity performance, with CTAB-GAN outperforming CTGAN. CTAB-GAN+ [?] is an extension of CTAB-GAN, which addresses privacy concerns by training a discriminator using differential private-SGD [?]. Importantly, CTAB-GAN+ demonstrates high fidelity performance.
VAE-based models. TVAE [?], proposed along with CTGAN, was designed to check whether CTGAN’s generator’s lack of access to the real data during training makes it weaker compared to models that do use these data for training their generator, such as models based on Variational Autoencoders (VAEs) [?]. Indeed, TVAE’s architecture showed an improvement over CTGAN in terms of fidelity. However, like CTAB-GAN+, the authors also considered the impact of their proposed models w.r.t. privacy and noted that TVAE’s access to the training data might make it unsuitable in downstream tasks where sensitive data are present.
Normalising flows-based models. Neural Spline Flows: Rather than using the common transformations found in normalising flows, ? [?] proposed using a fully-differentiable module based on monotonic rational-quadratic piecewise functions. This adaptation, combined with the inherent properties of normalising flows, which model the data as the output of an invertible differentiable transformation of a noisy sample drawn from a known distribution, can help synthesise high-fidelity tabular data. The authors note that such a result is conditional on having enough training data compared to the datasets dimensionality.
Diffusion-based models. FinDiff: Motivated by the need to preserve the statistical properties of real data of mixed types in real-world financial contexts, ? [?] proposes FinDiff, which is capable of generating mixed-type tabular financial data using a diffusion model to capture high-dimensional dependencies. Unlike another popular diffusion model, TabDDPM [?], which uses one-hot encoding for categorical features, FinDiff employs an embedding encoding for better handling of mixed data types. This leads to higher fidelity performance compared to the baselines, indicating that FinDiff effectively captures both feature-wise and joint distributions of the data.
Diffusion and AE-based models. AutoDiff: Proposed in [?], AutoDiff is a diffusion-based model designed to preserve statistical fidelity by modelling the joint distribution via an autoencoder, rather than modelling the distributions of individual features separately, as seen in GAN- and diffusion-based models like CTGAN and TabDDPM. More precisely, AutoDiff (i) uses the autoencoder to learn the distribution of the features (which can be of mixed types) in a continuous space, and then (ii) passes these continuous representations to the diffusion model, which generates latent representations that are then decoded back into representations in the original feature space. Moreover, like CTAB-GAN, it introduces an approach to handle mixed-type features by adding a new feature to the autoencoder that encodes the frequency of recurring values in a mixed-type feature.
Diffusion and VAE-based models. TabSyn: Introduced by ? [?], TabSyn trains a score-based diffusion model, like STaSy, but in a joint VAE-learned space of numerical and categorical features. Notably, TabSyn significantly reduces the runtime of diffusion-based models while outperforming baselines w.r.t. fidelity and utility.
6 Privacy
Synthetic data are particularly valuable in privacy-sensitive domains like finance and healthcare. So, there has been growing focus on generating data without revealing sensitive information that could identify entities from the original dataset.
6.1 Evaluation Protocol and Metrics
Privacy is typically evaluated by assessing the risk of attacks succeeding in targeting sensitive information. The most common attacks are: (i) membership attacks, which use ML classifiers or black-box attacks (e.g., see [?]) to determine whether an entity was part of the set used to train a model, (ii) attribute disclosure attacks, which try to gain access to sensitive attributes of an entity from the real dataset, and (iii) re-identification attacks, which try to map a synthetic data point back to the original dataset. To measure the risk of the attacks being successful, for (i), AUCROC is reported for the trained attack models (typically one for each class of a feature of interest), for (ii), precision and sensitivity (i.e., recall) are measured while varying the number of attributes known to the attacker, and, for (iii), common metrics include: the distance to the closest record (DCR) [?] (computed as the minimum L2 distance from a synthetic data point to each real data point), the nearest neighbour distance ratio (computed as the ratio between the distances to the closest and second-closest real neighbours for a synthetic record), but also k-anonymisation, delta presence, and the identifiability score, as seen in [?].
Finally, methods using differential privacy [?] introduce noise during training to reduce retention of original data features. Privacy is then evaluated using different privacy budget values, which determine the amount of added noise and balance utility and privacy (e.g., see [?]).
6.2 Methods
AE and GAN-based models. medGAN: Proposed by ? [?] for synthesising healthcare datasets, medGAN is one of the pioneering models for synthesising tabular data using deep generative models. medGAN generates distributed representations of health records using the generator of a GAN-based model, then decodes these representations using an autoencoder pretrained on the real data, and passes the decoded representations to the discriminator component of the GAN, which is trained to distinguish the synthetic from the real data. In terms of privacy, medGAN showed low attribute disclosure risks, except when an attacker focused on a small number of data points. IT-GAN: Similar to OCT-GAN but designed to enhance privacy robustness, IT-GAN [?] employs a generator based on neural ODEs to balance utility and privacy protection. Instead of directly synthesising samples in the input space, the generator produces hidden representations, as seen also in medGAN. Notably, it ensures that the hidden representation space matches the input dimensionality, enabling the application of ODEs.
GAN-based models. TableGAN: Another model that marked an important milestone for deep generative modelling of tabular data is TableGAN [?], which, unlike medGAN, can model continuous features. While it is based on a standard GAN model, consisting of a generator and a discriminator, TableGAN has an additional component: a classifier that predicts the value of the target feature for synthetic samples using semantic integrity constraints learned from the real data. As exemplified in the original paper, a sample with values of $60$ and $1$ for features Cholesterol and Diabetes, respectively, does not align with the background knowledge, as the cholesterol level is too low for the target feature to indicate a diabetes diagnosis. Focusing on ensuring privacy, the authors argue for the use of synthesiser models, which do not learn bijective relationships between real and fake data, making them less vulnerable to re-identification attacks, unlike prior perturbation or anonymisation methods. Indeed, TableGAN shows that it can synthesise tabular data with a low risk of re-identification, membership and attribute disclosure attacks to succeed. PATE-GAN: Another model that is based on a GAN architecture, but alters it for preserving privacy, is PATE-GAN [?]. Rather than using the standard GAN discriminator, PATE-GAN relies on a modified Private Aggregation of Teacher Ensemble (PATE) [?] mechanism to ensure that the discriminator is differentially private. More specifically, a number of teacher-discriminators are created and trained to improve their loss w.r.t. the generator, which adjusts its loss according to the single student-discriminator in the mechanism. In turn, the student-discriminator adjusts its loss according to the teacher-discriminators and allows for backpropagation to the generator, closing the loop. Critically to the privacy requirement, specialising each of the teacher-discriminators on small partitions of the training set ensures that the effect of individual samples is small and, thus, higher differential privacy. CuTS: ? [?] proposed CuTS as a customisable framework, which can adapt to user-defined specifications, such as background knowledge constraints, but also to differentially private training. In the private setting, CuTS builds on the DP iterative framework from [?], but uses its own GAN-based generator, which is equipped with a mechanism that matches the generator’s output distribution to the real data distribution by comparing marginals. Additionally, CuTS can incorporate background knowledge in the form of propositional logical constraints that each data point must satisfy via a new loss term. Further, rejection sampling is used to ensure that the model’s outputs satisfy the constraints. Although the framework works only with discrete features, it takes a step towards synthesisers for tabular data that account for multiple requirements, including alignment and privacy.
7 Other models
In (§3-6), we compared deep generative models originally proposed for tabular data synthesis. Here, we discuss seminal models first developed in other fields and later adapted for tabular data. WGAN: Originally designed for computer vision tasks, WGAN [?] was introduced in a paper that analyses ways to measure the distance between the real and model distributions, arguing that weaker loss functions induce weaker model topologies, thus weak models for the real distribution. The authors compare the Wasserstein distance with traditional probability distance measures, including Kullback-Leibler (KL) divergence, Jensen-Shannon divergence, and total variation distance, showing that the Wasserstein distance has properties that would be better suited for learning distributions in low dimensional manifolds. They thus propose GAN models that use the Wasserstein distance as loss function. WGAN-GP: A popular extension of WGAN is WGAN-GP [?], which added a new term in the loss that penalised the gradient norm of the discriminator w.r.t. the discriminator’s input, providing an alternative to the weight clipping that was leading to the generator retaining the real data points and adding Gaussian noise to fill the gaps. VEEGAN: VEEGAN [?] jointly trains a generator and a novel reconstructor network in order to avoid mode collapse. The reconstructor learns to map the real data distribution to Gaussian random noise, approximating the inverse of the generator and encouraging the generator to map the noise distribution back to the real data distribution.
Figure 2: Visualisation of the surveyed models based on the requirements they address.
8 Discussion and Future Directions
In this survey, we identified four key requirements for the effective deployment of synthetic data, along with the relevant evaluation procedures and deep generative models, at the same time categorising the latter by their model architecture.
Related work. Previous surveys generally compare methods by the architecture-specific properties of the models. For example, both ? (?) and ? (?) group methods mainly by their architectures, while also describing the necessary data transformation and regularisation techniques. [?], provide an overview of the models proposed in the field and the metrics used to evaluate them, while not taking this requirement-centric perspective which helps both newcomers and seasoned practitioners and researchers in the field to find the perfect model and evaluation protocol for their needs. Further, they exclude diffusion-based and LLM-based architectures.
Future Directions. Tabular data generation is a rapidly evolving field still in its early stages. As illustrated in Figure 2, many early approaches have prioritised individual objectives—typically utility or privacy—leading to highly specialised models that lack broad applicability. This fragmentation highlights a key challenge: the need for models that can balance multiple, often competing, requirements.
It is worth to be noted that alignment has historically received limited attention but is now emerging as a crucial factor for generating realistic data. With the increasing deployment of generative models in real-world applications and the growing prominence of neurosymbolic AI and safe AI [??], we anticipate that alignment will soon become a fundamental requirement in the field.
Moreover, Figure 2 reveals a striking gap: no existing model successfully integrates all the core requirements we consider. As the field matures, we expect new requirements will emerge, and the hybridisation of techniques will drive the development of more versatile models. Bridging these gaps will be essential for advancing the field and enabling broader, more reliable applications of tabular data generation.
Acknowledgments
Mihaela Cătălina Stoian is supported by the EPSRC under the grant EP/T517811/1.